Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 16:00 Monday 17 May 2021 (51813) Print this report
Operator Report - Morning shift

The planned commissioning activity started around 7:00 UTC.

Unfurtunately the ISYS was not collaborative as we suffered of many fast unlocks (see attached plot) which prevented the smooth running of the activity.

At around 9:00 UTC we contacted the ISYS on-call, see On-call intervention.

 

Other activities:

- TCS: CO2 guardian system ON, all the related flags have been unshelved on DMS under request of Ilaria;

- SVS: Optimization of mode matching on EQB1 for SQZ measurement (Sorrentino, Garaventa)

- SIN: PSD installation on SQB2 (Polini, Vardaro, ?)

ISYS
(17-05-2021 09:00 - 17-05-2021 10:45) On site
Status: Ended
Description: IMC unlocks
Actions undertaken: see entry 51810

Images attached to this report
Search Help
×

Warning

×